Pls tell i am not understanding this equation
 
Sin(A+A2+An)=  

Grade:11

1 Answers

Aditya Gupta
3 years ago
the ques isnt clearly written so kindly upload a photo.
if you want to know how to find a formula for sin(x+y+z), then let u=y+z
so sin(x+y+z)= sin(x+u)= sinx.cosu+sinu.cosx= sinx(cosycosz – sinysinz) + cosx(sinycosz+sinzcosy)
= sinxcosycosz+sinycosxcosz+sinzcosxcosy – sinxsinysinz
KINDLY APPROVE :))